OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (−1, −2) and (3, 10). y+____=____(x+1)

OpenStudy (anonymous):

Correction- (-1,-2) and (3,10)

OpenStudy (anonymous):

@dan815 @cwrw238

OpenStudy (anonymous):

@ganeshie8 @paki

OpenStudy (anonymous):

@satellite73 @TheSmartOne @mathmate @mathmath333 @mathstudent55

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{(x_1,y_1)=(-1,-2)\\ (x_2,y_2)=(3,10)\\ y-y_1=m(x-x_1)\\~\\~\\ m=\dfrac{y_1-y_2}{x_1-x_2}\\ =\dfrac{-2-10}{-1-3}\\ =\dfrac{-12}{-4}\\ =3\\~\\~\\ y-(-2)=3(x-(-1))\\ y+2=3(x+1)\\ }\end{align}\)